for solve for y
1. y-3x=7
2. 2x+y=5
solve for a
a+3b=-2

Answer by Pieter(35)

for y: just put everything that is not y to the other side.
1)y - 3x = 7
or y = 3x + 7
2) 2x + y = 5
or y = 5 - 2x
for a: just put everything that is not a to the other side
a + 3b = -2
or a = -2 - 3b
Greets
Pieter Gurdebeke